a. the temperature of 200g of water increases from 20 0C to 50 0C in 10 minutes. Calculate:

i. the amount of heat gained by the water

ii. the average rate at which the water is gaining heat


(specific heat capacity of water = 4200 J kg-1 0C-1)


b. Determine the amount of heat that must be removed from 1 kg of water at 40 0C to convert it to ice at -5 0C.

(Obtain any relevant data that you need from the internet. Cite the source of that data in your answer)


1
Expert's answer
2022-01-31T14:23:26-0500

a)

(i) the amount of heat gained by the water

"Q=cm\Delta t=4200*0.2*(50-20)=25200\:\rm J"

(ii) the average rate at which the water is gaining heat

"P=\frac{Q}{\tau}=\frac{25200}{10*60}=42\:\rm W"

b)

"Q=c_wm\Delta t_1+\lambda m+c_{ice}m\Delta t_2\\
=4200*1*40+330000*1+2100*1*5\\
=508500\:\rm J"